OpenStudy (amypurplediaz28):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) −1 −9 0 −1 1 7 g(x) g(x) = 3x − 2 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

jimthompson5910 (jim_thompson5910):

`Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points)` focus on the first two rows of the table x y -1 -9 0 -1 hopefully you agree that those two rows mean we have these two points (-1,-9) and (0,-1) ?

OpenStudy (amypurplediaz28):

yea i do

jimthompson5910 (jim_thompson5910):

what is the slope of the line through the two points (-1,-9) and (0,-1) ?

jimthompson5910 (jim_thompson5910):

slope formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (amypurplediaz28):

ok hold on

OpenStudy (amypurplediaz28):

-8/-1

jimthompson5910 (jim_thompson5910):

which reduces to what?

OpenStudy (amypurplediaz28):

idk

jimthompson5910 (jim_thompson5910):

two negatives divide to make a positive, so 8/1 = 8

OpenStudy (amypurplediaz28):

i think that is all it reduces to

jimthompson5910 (jim_thompson5910):

(-8)/(-1) = 8/1 = 8

OpenStudy (amypurplediaz28):

oh that make sense

jimthompson5910 (jim_thompson5910):

so the slope of f(x) is 8

jimthompson5910 (jim_thompson5910):

what is the slope of g(x) ?

OpenStudy (amypurplediaz28):

how do u find that

jimthompson5910 (jim_thompson5910):

are you familiar with y = mx+b?

OpenStudy (amypurplediaz28):

yes

jimthompson5910 (jim_thompson5910):

with y = mx+b, m is the slope

jimthompson5910 (jim_thompson5910):

so compare y = 3x-2 with y = mx+b do you notice how the '3' and the 'm' match up? do you see how they're in front of the x?

OpenStudy (amypurplediaz28):

yea

jimthompson5910 (jim_thompson5910):

so you would agree that the slope of g(x) is 3 ?

OpenStudy (amypurplediaz28):

yes i agree

jimthompson5910 (jim_thompson5910):

summary: slope of f(x) is 8 slope of g(x) is 3 we can see that f(x) has a larger slope

OpenStudy (amypurplediaz28):

yes

OpenStudy (amypurplediaz28):

so that is part b right

OpenStudy (amypurplediaz28):

also what would i write as a comparison in part a

OpenStudy (amypurplediaz28):

@jim_thompson5910

OpenStudy (amypurplediaz28):

@jim_thompson5910

OpenStudy (amypurplediaz28):

are u there

jimthompson5910 (jim_thompson5910):

sorry about that

jimthompson5910 (jim_thompson5910):

that was all for part A

OpenStudy (amypurplediaz28):

so we begin with part B

jimthompson5910 (jim_thompson5910):

yes let's look at part B now

jimthompson5910 (jim_thompson5910):

`Part B: Which function has a greater y-intercept? Justify your answer. (4 points)` Look at the table. What is the y value when x = 0?

OpenStudy (amypurplediaz28):

-1

jimthompson5910 (jim_thompson5910):

so -1 is the y intercept for f(x)

jimthompson5910 (jim_thompson5910):

what is the y intercept for g(x) ?

jimthompson5910 (jim_thompson5910):

hint: y = mx+b m = slope b = y intercept

OpenStudy (amypurplediaz28):

i am not sure

jimthompson5910 (jim_thompson5910):

compare y = 3x-2 to y = mx+b m = 3 b = -2 agreed?

OpenStudy (amypurplediaz28):

agreed

jimthompson5910 (jim_thompson5910):

b = -2 is the y intercept of g(x)

jimthompson5910 (jim_thompson5910):

y intercept of f(x) is -1 y intercept of g(x) is -2

jimthompson5910 (jim_thompson5910):

so we see that f(x) has the greater y intercept because -1 is larger than -2

OpenStudy (amypurplediaz28):

so g(x) has the higher y intercept

jimthompson5910 (jim_thompson5910):

-2 is below -1

jimthompson5910 (jim_thompson5910):

draw out a vertical number line like so 

OpenStudy (amypurplediaz28):

oh right so -1 is higher i was thinkin that they were positive sorry

jimthompson5910 (jim_thompson5910):